Penny Company manufactures only one product and uses a standard cost system. The following information is from Penny’s records for May:

Direct labor rate variance $15,000 favorable
Direct labor time variance $31,200 unfavorable
Standard hours per unit produced 2.00
Standard rate per hour $26

During May, the company used 12.50% more hours than the standard allowed.

A. What were the total standard hours allowed for the units manufactured during the month?

standard hours

B. What were the actual hours worked?

actual hours

C. How many actual units were produced during May?

units produced
Answer #1

A. Direct labor time variance = (Standard hours -actual hours)* standard rate

-31200 = (Standard hours - Standard hours*1.1250)*26

Standard hours = 9600 hours

B. Actual hours worked =  Standard hours*1.1250 = 9600*1.1250 = 10800 hours

C. Actual units produced = 9600 / 2 = 4800 units